# Ivy-Fake: A Unified Explainable Framework and Benchmark for Image and Video AIGC Detection

### v1

@@ 3.4 Comparison with Existing Datasets
FakeBench (Li et al., 2024c) - 6 Image 3K 3K
VANE-Bench (Bharadwaj et al., 2024) 101 5 Image* 2K 1K
LOKI (Ye et al., 2025) 99 $\sim$16 Image+Video 3K 0
FakeClue (Ye et al., 2025) 120 26 Image 68K 36K
Ivy-Fake 530 >30 Image+Video 74K 86K
In contrast, datasets such as FakeClue (Wen et al., 2025) and FakeBench (Li et al., 2024c) provide only limited annotated samples, thereby restricting their applicability primarily to evaluation scenarios rather than serving as robust resources for model training. Similarly, LOKI (Ye et al., 2025), while offering explainability-focused annotations, is confined to a single modality (images) and evaluates across fewer dimensions. This constraint limits its suitability for comprehensive research on multimodal AIGC detection and explainability. By offering large-scale, richly annotated datasets spanning images and videos, Ivy-Fake  establishes a more holistic and versatile benchmark for advancing explainable AIGC detection in multimodal contexts.
@@ 4 Methodology
This study targets at an unified benchmark of AIGC detection for both image and video content. Our preliminary investigations revealed that existing MLLMs exhibit inadequate performance on these tasks. To overcome this limitation, we propose Ivy-xDetector, a multimodal large language model designed explicitly for robust and explainable AIGC detection. The following subsections detail our instructional dataset construction and the progressive multimodal training framework employed in developing Ivy-Detector.
@@
Figure 3: Overview of the three-stage training pipeline for Ivy-Detector, including general video understanding, detection instruction tuning, and interpretability instruction tuning.
4.1 Ivy-Detector model
We introduce Ivy-xDetector, an architecture that adheres to the LLaVA (Liu et al., 2023; Li et al., 2024a) paradigm and is initialized using Ivy-VL-LLaVA (Zhang et al., 2024) weights. As illustrated in Figure 3, The Ivy-xDetector framework comprises three core components: a Visual Encoder, a Visual Projector, and a Large Language Model (LLM). We employ SigLIP (Zhai et al., 2023) as the visual backbone to process input images and frames extracted from video sequences. To enable fine-grained detection, particularly for high-resolution images, we implement a dynamic resolution strategy. Input images are partitioned into multiple $384\times 384$ sub-images, which are then collectively fed into the Visual Encoder. This approach supports an effective input resolution of up to $2304\times 2304$. For video inputs, individual frames are resized to $384\times 384$. Post-encoding by the Visual Encoder, the resulting features undergo a pooling operation, compressing the token sequence to one-quarter of its original length. Crucially, to preserve rich temporal information from video data, we eschew temporal compression of video features. Features from all frames are concatenated and subsequently processed by the LLM. Through joint training on image and video datasets, Ivy-xDetector acquires a unified understanding of both modalities, enabling it to perform explainable detection tasks for both images and videos.
4.2 Progressive Multimodal Training Framework
To effectively train Ivy-xDetector and fully leverage our curated instruction dataset, we devise a progressive multi-modal training framework. This framework employs a staged optimization strategy to incrementally enhance the model’s proficiency in both AI-Generated Content (AIGC) detection and the generation of associated explanations.
Stage 1: Imbuing Video Understanding Capabilities.
We initialize Ivy-xDetector using the Ivy-VL-LLaVA model (sub-4B parameter scale), which has demonstrated state-of-the-art (SOTA) performance on image-text benchmarks but lacks prior exposure to video data. To address this, we curate a dataset of 3 million video-text pairs, aggregated from sources such as VideoChatFlash (Li et al., 2024b) and VideoLLaMA3 (Zhang et al., 2025), to equip the model with fundamental video comprehension abilities. Further details regarding these datasets are provided in Appendix.
Stage 2: AIGC Detection Fine-tuning.
Previous iterations of the base model were primarily trained on authentic (non-generated) datasets. To specialize Ivy-Fake for AIGC detection, we compile a targeted dataset for instruction fine-tuning, drawing from established datasets including Demamba, FakeClue, and WildFake. This dataset undergoes rigorous preprocessing, including the removal of corrupted or low-quality samples. For video data, frames are extracted at a uniform rate of 1 frame per second (fps). The core objective of this stage is to train the Multi-modal Large Language Model (MLLM) component for binary AIGC discrimination—classifying content as either “real” or “fake.” Given the focused nature of this binary classification task, a single, consistent instruction template is utilized. This stage aims to cultivate high-accuracy classification capabilities across diverse sources and types of AIGC.
Stage 3: Joint Optimization for Instruction-Driven Detection and Explainability.
This culminating stage of training is designed to endow Ivy-xDetector with the capacity to generate high-quality, human-understandable explanations while preserving the AIGC detection accuracy achieved in Stage 2. Our preliminary experiments revealed that sequentially fine-tuning for explainability after the detection-focused training (Stage 2) often resulted in the model struggling to adhere to detection-specific instructions. Consequently, in Stage 3, we adopt a joint training approach, concurrently fine-tuning the model on a combined dataset comprising the AIGC detection data from Stage 2 and newly introduced explainability-focused instruction data. The instructions in this phase are crafted to elicit detailed, step-by-step reasoning from the model, thereby placing greater demands on its generative and inferential capabilities. The model emerging from this stage is expected to not only accurately perform AIGC detection but also provide comprehensive and persuasive rationales for its classifications.
Through this three-stage progressive training paradigm, we aim for Ivy-xDetector to systematically develop a comprehensive skill set: from discerning subtle AIGC artifacts and making accurate classifications, to articulating coherent and justifiable explanations. This holistic approach is designed to ensure Ivy-xDetector exhibits robust and superior performance on complex AIGC detection challenges.
@@ 5 Experiments
We perform extensive experiments to assess both detection and explanation capabilities. In particular, the proposed method is evaluated on the classification (real/fake) tasks and reasoning tasks for both image and video content using the proposed unified framework. For the classification task, we test our model on both image and video content to detect the synthesic content. We report standard accuracy (Acc) and macro-averaged F1 score (F1) to assess the model’s ability to distinguish real from fake instances. For the reasoning task, we measure the similarity between the model’s reasoning process and the reference annotations using the ROUGE-L score (Lin, 2004), which captures the longest common subsequence between predicted and reference texts, reflecting token-level overlap. Since ROUGE-L may fail to fully capture the fidelity of reasoning steps, we adopt an LLM-as-a-judge evaluation paradigm (Zheng et al., 2023b), following the FakeBench protocol (Li et al., 2024c), which assesses model responses along four dimensions: (1) Completeness: It reflects the extent to which the response fully addresses all aspects of the user’s question. More complete responses should incorporate information aligning well with the “golden clues” or reference answers. Incomplete or partially answered responses will receive lower scores. (2) Relevance: Measure how closely the content relates to the original annotation; (3) Level of Detail: Assess whether the response includes enough examples or elaborations; (4) Explanation: Verify the accuracy and consistency of explanations for any causes mentioned. Each response is scored using GPT-4o mini (Achiam et al., 2023) under a unified evaluation prompt (see Appendix 2), which instructs the model to act as an impartial judge and assign a score from 1 to 5. To ensure consistency and reduce variance, each response is rated over five independent rounds, and the final score is obtained by averaging these outputs.
@@
Table 2: Comparison on the Genimage (Zhu et al., 2023b). Accuracy (%) of different detectors (rows) in detecting real and fake images from different generators (columns). The best result and the second-best result are marked in bold and underline, respectively.
@@
CNNSpot (Wang et al., 2020) 52.80 96.30 95.90 50.10 39.80 78.60 53.40 46.80 64.21
F3Net (Qian et al., 2020) 50.10 99.90 99.90 49.90 50.00 99.90 49.90 49.90 68.69
DIRE (Wang et al., 2023) 60.20 99.90 99.80 50.90 55.00 99.20 50.10 50.20 70.66
GenDet (Zhu et al., 2023a) 89.60 96.10 96.10 58.00 78.40 92.80 66.50 75.00 81.56
PatchCraft (Zhong et al., 2023) 79.00 89.50 89.30 77.30 78.40 89.30 83.70 72.40 82.30
AIDE (Yan et al., 2025) 79.38 99.74 99.76 78.54 91.82 98.65 80.26 66.89 86.88
Ivy-Det 92.37 99.20 99.23 99.35 99.33 99.08 99.20 99.16 98.36
Ivy-xDet 90.26 98.28 98.09 98.58 98.38 98.20 98.29 98.23 97.29
5.1 Image Content Classification
We evaluated our image AIGC detector on the GenImage (Zhu et al., 2023b) and Chameleon (Yan et al., 2025) benchmarks. GenImage comprises seven subsets generated by leading models, i.e., Midjourney, Stable Diffusion v1.4 & v1.5, ADM, GLIDE, Wukong, VQDM, and BigGAN. We compared against five state-of-the-art detectors, which are CCNNSpot (Wang et al., 2020), F3Net (Qian et al., 2020), DIRE (Wang et al., 2023), GenDet (Zhu et al., 2023a), PatchCraft (Zhong et al., 2023), and AIDE (Yan et al., 2025). As Table 2 shows, our method raises mean accuracy from 86.88 % to 98.36 %. For the subsets like BigGAN, we can even improve 32.27%, which shows the superiority of our new benchmark.
On Chameleon, we compared with ten detection methods, i.e., CNNSpot, F3Net, DIRE, GenDet, PatchCraft, and AIDE. Following work (Yan et al., 2025), both the overall accuracy and the separate accuracy on synthetic content (fake) and real data (real) are reported. As shown in the Table 3, our methods (last two columns), at least improved 20% when compared with any of previous best methods. This further demonstrate the superiority of our method on the image level AGIC detection.
Table 3: Comparison on the Chameleon (Yan et al., 2025). Accuracy (%) of different detectors (rows) in detecting real and fake images. For each training dataset, the first row is overall accuracy, the second row is “fake/real” class accuracy.
CNNSpot FreDect Fusing GramNet LNP UnivFD DIRE PatchCraft NPR AIDE Ivy-Det Ivy-xDet
60.89 57.22 57.09 60.95 58.52 60.42 59.71 56.32 58.13 65.77 85.20 83.39
9.86/99.25 0.89/99.55 0.02/99.98 4.76/99.66 7.72/96.70 85.52/41.56 11.86/95.67 3.07/96.35 2.43/100.00 26.80/95.06 76.21/95.61 74.49/93.71
5.2 Video Content Classification
We further evaluate our unified model on video-level AIGC detection using the GenVideo dataset (Chen et al., 2024a), the largest benchmark for generated video detection, in a many-to-many generalization setting. Without modification, we apply the same architecture used for image analysis and compare against four state-of-the-art methods: F3Net (Qian et al., 2020), NPR (Tan et al., 2024), STIL (Gu et al., 2021), and DeMamba-XCLIP-FT (Chen et al., 2024a). Performance is reported in terms of recall (R), F1 score (F1), and average precision (AP).
Figure 4: Showcase of the Video artifact detection.
Figure 4 shows one case for the video detection. Our model progressively conducts the spatial and temporal analysis to better capture generation content. As Table 4 shows, our model outperforms all baselines, achieving over 99 % accuracy across most generative sources. Notably, on the most challenging “HotShot” subset, we achieve a recall of 99.57 %, versus 65.43 % for the previous best method, underscoring our approach’s superior effectiveness in video-level AIGC detection.
5.3 Image and Video generation content Reasoning
This section evaluates MLLMs’ reasoning over image- and video-generated content, in addition to their detection performance across both modalities. We benchmark our model against four leading LMMs on the full Ivy-Fake: two open-source (Qwen2.5-7B (Bai et al., 2025) and InternVL2.5-8B (Chen et al., 2024b, c)) and two proprietary (GPT-4V (Achiam et al., 2023) and Gemini 2.5 Pro (Team et al., 2023)). Detailed reasoning results are provided in the Appendix. Our approach not only achieves superior accuracy but also delivers more transparent explanations than any baseline. Notably, unlike methods such as LOKI (Ye et al., 2025), which express interpretability via artifact coordinates or heatmaps, our model generates natural-language descriptions of visual artifacts that are immediately accessible to human readers.
Table 4: Comparisons to the GenVideo. F1 score (F1), recall score (R) and average precision (AP) on the many-to-many generalization task. “Demamba-XCLIP-FT” is abbreviated as “Demamba”.
@@ 5.3 Image and Video generation content Reasoning
AP 0.9332 1.0000 0.9997 0.8555 0.9897 0.9960 0.9998 1.0000 0.9777 0.9575 0.9710
Ivy-Det R 1.0000 0.9986 0.9993 0.9957 0.9943 0.9943 1.0000 1.0000 0.9986 0.9601 0.9948
F1 1.0000 0.9993 0.9996 0.9979 0.9971 0.9971 1.0000 1.0000 0.9240 0.9796 0.9974
AP 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
Ivy-xDet R 0.9821 1.0000 0.9935 0.9771 0.9943 0.9900 1.0000 0.9993 0.9871 0.8690 0.9839
F1 0.9910 1.0000 0.9967 0.9884 0.9971 0.9950 1.0000 0.9996 0.9935 0.9299 0.9919
AP 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000

### v2

@@ 2.3 Comparison with Existing Datasets
FakeBench [liu2024mmfakebench] - 6 Image 3K 3K
VANE-Bench [bharadwaj2024vanebench] 101 5 Image* 2K 1K
LOKI [loki_ye2024loki] 226.5 $\sim$16 Image+Video 3K 0
FakeClue [fakeclue_wen2025spot] 119.6 26 Image 68K 36K
Ivy-Fake 451.7 >30 Image+Video 53K 78K
In contrast, datasets such as FakeClue [fakeclue_wen2025spot] and FakeBench [liu2024mmfakebench] provide only limited annotated samples, thereby restricting their applicability primarily to evaluation scenarios rather than serving as robust resources for model training. Similarly, LOKI [loki_ye2024loki], while offering explainability-focused annotations, is confined to a single modality (images) and evaluates across fewer dimensions. This constraint limits its suitability for comprehensive research on multimodal AIGC detection and explainability.
We identified a long-tail distribution in the distilled dataset, with a small portion of samples exhibiting token lengths concentrated at both extremes. To mitigate the influence of these outliers, we retained only the central 90% of the distribution, filtering out the top and bottom 5% of samples, thereby keeping data with token lengths between 60 and 800. As shown in Figure 2 panel (a), Ivy-Fake exhibits the longest token length distribution, with an average token length of 451.7. Additionally, to conduct a fine-grained analysis of the semantic dimensions embedded in the explanatory language across different interpretability datasets, we employ Qwen3-32B [qwen3] for dimension extraction. To enhance the accuracy of the extraction, we embed supplementary contextual information from each dataset into the model input. For example, in LOKI [loki_ye2024loki], the bounding box coordinates are wrapped with <bbox></bbox> tags, while metadata such as coordinates, titles, and fine-grained descriptions are also included as input to the model. Moreover, we set the Temperature to 0.8 and perform three extraction passes for each sample, retaining only the dimensions consistently identified across all runs. The final aggregated results are visualized in panel (b).
Figure 3: Comparison between Ivy-Fake and FakeVLM [fakeclue_wen2025spot] (NeurIPS 2025). The Ivy-Fake dataset provides richer and more fine-grained interpretability dimensions.
As shown in Figure 3, Ivy-Fake outperforms FakeVLM in terms of interpretability. While FakeVLM only explains superficial factors such as lighting and texture, Ivy-Fake considers multiple semantic dimensions, including hair, skin, and background, leading to a more comprehensive understanding of forged content. The more image and video cases can be found in Appendix 14.
2.4 Quality Control
To mitigate the impact of hallucinations from Gemini 2.5 Pro during question generation, we sampled 8,000 cases for manual review. Each chain-of-thought (CoT) involving Gemini underwent at least two rounds of human verification. A total of ten users, including senior university students and regular participants, contributed to the verification process, which required approximately 1,000 hours in total. Each case was tested by at least two users to ensure robustness across different modalities of synthetic data. After this rigorous filtering, we retained 5,000 cases as the benchmark test set, consisting of 2,500 image samples and 2,500 video samples, with each modality containing 1,250 real and 1,250 fake instances. The complete dataset distribution is provided in Appendix 10.
@@ 3 Methodology
Our preliminary investigations revealed that existing MLLMs exhibit inadequate performance on these tasks. To overcome this limitation, we propose Ivy-xDetector, a multimodal large language model designed explicitly for robust and explainable AIGC detection. The overall training pipeline is illustrated in Figure 4.
@@
Figure 4: Overview of the three-stage training pipeline for Ivy-xDetector, including general video understanding, detection instruction tuning, and interpretability instruction tuning.
Stage 1: Cold Start for Instruction-Driven Detection and Explainability.
As shown in Figure 3, due to limited reasoning ability, existing models struggle to detect AIGC content and generate reliable explanations. To address this, we inject Gemini-generated detection and explanatory COT into the training process, thereby improving the model’s fine-grained artifact perception and explanation quality. This stage aims to initialize Ivy-xDetector with instruction-following and explainability capabilities. The resulting model not only performs accurate AIGC detection but also generates coherent, human-understandable rationales for its classifications.
Table 3: Performance comparison of models on image and video tasks. “Auto Metrics” include Acc, F1. “GPT Assisted” includes four subjective criteria: Comprehensiveness, Relevance, Detail, and Explanation. Bold indicates the best result, and underline indicates the second best.
Model Image Video Overall
Auto Metrics GPT Assisted Auto Metrics GPT Assisted Auto Metrics GPT Assisted
Acc/F1/ROUGE-L/SIM Com./Rel./Det./Exp. Acc/F1/ROUGE-L/SIM Com./Rel./Det./Exp. Acc/F1/ROUGE-L/SIM Com./Rel./Det./Exp.
Closed-source MLLMs
GPT-4o 0.725/0.723/0.108/0.525 2.34/3.20/2.04/3.26 0.448/0.579/0.072/0.451 1.79/2.35/1.67/2.40 0.587/0.663/0.090/0.488 2.07/2.78/1.85/2.83
Gemini-2.5-Flash 0.747/0.737/0.263/0.733 3.94/4.11/4.04/4.09 0.810/0.811/0.246/0.723 4.00/4.37/4.03/4.36 0.779/0.776/0.254/0.728 3.97/4.24/4.04/4.22
Open-source MLLMs
7B-Parameters Models
InternVL3.5-8B 0.605/0.602/0.194/0.680 2.83/3.49/2.69/3.32 0.574/0.588/0.188/0.664 2.75/3.35/2.68/3.28 0.589/0.596/0.191/0.672 2.79/3.42/2.69/3.30
MiMo-VL-7B 0.662/0.637/0.121/0.593 1.99/2.80/1.85/2.89 0.778/0.783/0.112/0.580 2.04/2.91/1.90/3.19 0.720/0.715/0.116/0.586 2.01/2.86/1.87/3.04
Qwen2.5-VL-7B 0.013/0.026/0.006/0.264 1.02/1.03/1.01/1.52 0.092/0.159/0.015/0.280 1.14/1.23/1.12/2.21 0.053/0.096/0.010/0.272 1.08/1.13/1.07/1.86
LLaVA-OneVision-1.5-8B 0.500/0.333/0.080/0.499 1.51/2.62/1.49/2.38 0.500/0.333/0.068/0.481 1.49/2.26/1.37/2.19 0.500/0.333/0.074/0.490 1.50/2.44/1.43/2.28
MiniCPM-V-4.5 0.666/0.680/0.169/0.637 3.20/3.93/3.06/3.60 0.491/0.505/0.152/0.627 2.83/3.66/2.76/3.36 0.579/0.610/0.161/0.632 3.01/3.80/2.91/3.48
3B-Parameters Models
Qwen2.5-VL-3B 0.641/0.612/0.023/0.391 1.19/1.33/1.18/3.28 0.689/0.686/0.017/0.381 1.42/1.56/1.40/3.68 0.665/0.652/0.020/0.386 1.31/1.45/1.29/3.48
Gemma-3-4b-it 0.408/0.477/0.170/0.576 2.55/3.36/2.46/3.11 0.396/0.482/0.149/0.561 2.30/3.03/2.37/2.95 0.402/0.482/0.159/0.568 2.43/3.19/2.42/3.03
InternVL3.5-2B 0.602/0.573/0.177/0.648 2.62/3.29/2.51/3.13 0.435/0.459/0.159/0.631 2.46/3.16/2.42/2.99 0.518/0.518/0.168/0.640 2.54/3.22/2.47/3.06
InternVL3.5-4B 0.651/0.652/0.190/0.660 3.01/3.68/2.95/3.58 0.614/0.617/0.181/0.653 2.93/3.61/2.83/3.52 0.632/0.635/0.186/0.656 2.97/3.64/2.89/3.55
Ivy-xDetector 0.831/0.831/0.283/0.714 3.54/4.04/3.61/3.85 0.897/0.897/0.300/0.726 3.72/4.12/3.75/4.24 0.864/0.864/0.291/0.720 3.63/4.08/3.68/4.05
Stage 2: Sparse Rewards in Fine-grained Visual forgery Reasoning via Reinforcement Learning
Although the fine-tuned model from Stage 1 demonstrates improved artifact awareness and interpretability, its generalization ability remains limited. To further enhance the model’s capacity for producing consistent and human-understandable explanations, we adopt the Group Relative Policy Optimization (GRPO) algorithm [shao2024deepseekmathpushinglimitsmathematical]. GRPO enables reinforcement learning using only binary classification samples, achieving efficient and data-light optimization. A detailed description of GRPO and the reward functions for Ivy-xDetector can be found in Appendix 8.
We construct binary real/fake pairs from multiple datasets, each containing both authentic and synthetic images and videos. As illustrated in Figure 3, Ivy-xDetector takes the visual input and extracts the text enclosed within the <conclusion></conclusion> tags, where the predicted label is either real or fake. The quantitative results of different training stages are presented in Table 7.
@@ 4 Experiments
4.1 Experimental Details
Baselines
We primarily evaluate three closed-source models and three open-source models on our Ivy-Fake. The closed-source models are GPT-4o [gpt4_achiam2023gpt] and Gemini2.5 Flash [genimi_team2023gemini]. For the open-source models, we select models of comparable size: InternVL3.5 [chen2024internvl], Qwen2.5-VL [bai2025qwen2], MiMo-VL [coreteam2025mimovltechnicalreport], LLaVA-OneVision-1.5 [LLaVA-OneVision-1.5, lillava] and MiniCPM-V-4.5 [yao2024minicpm, yu2025minicpmv45cookingefficient]. Ivy-xDetector denotes our fully RL-trained model after the GRPO optimization stage.
Evaluation
@@ 4.1 Experimental Details
We report standard accuracy (Acc) and macro-averaged F1 score (F1) to assess the model’s ability to distinguish real from fake instances. For the reasoning task, we measure the similarity between the model’s reasoning process and the reference annotations using the ROUGE-L score [lin2004rouge] and BertScore [bert-score, he2021deberta], which captures the longest common subsequence between predicted and reference texts, reflecting token-level overlap. Since ROUGE-L may fail to fully capture the fidelity of reasoning steps, we adopt an LLM-as-a-judge evaluation paradigm [zheng2023judging], following the FakeBench protocol [liu2024mmfakebench], which assesses model responses along four dimensions: (1) Completeness: It reflects the extent to which the response fully addresses all aspects of the user’s question. More complete responses should incorporate information aligning well with the “golden clues” or reference answers. Incomplete or partially answered responses will receive lower scores. (2) Relevance: Measure how closely the content relates to the original annotation; (3) Level of Detail: Assess whether the response includes enough examples or elaborations; (4) Explanation: Verify the accuracy and consistency of explanations for any causes mentioned. Each response is scored using GPT-4o mini [gpt4_achiam2023gpt] under a unified evaluation prompt (see Appendix 12), which instructs the model to act as an impartial judge and assign a score from 1 to 5.
Figure 5: Ivy-xDetector performance on video detection. The model provides fine-grained spatiotemporal reasoning chains for explainable analysis.
Implementation Details
Our model was trained using the AdamW [adamw_Loshchilov2017FixingWD] optimizer with a cosine learning rate schedule. During the SFT stage, the model was trained for one epoch with a batch size of 1, which ensures stable convergence on large-scale multimodal data. In the subsequent RL stage, we adopted the GRPO algorithm, training for one epoch with the same batch size and a learning rate of $1\times 10^{-5}$. Within GRPO, the group size $n$ was set to 8, the warm-up ratio to 0.01, the temperature to 0.9. Training the SFT and RL stages required approximately 10 and 50 hours, respectively, on a system equipped with 32 A100 GPUs.
For video-based inputs, we configured the preprocessing pipeline to sample frames at a rate of 1 frame per second, with a maximum of 6 frames per clip to ensure temporal coherence while maintaining computational efficiency. Across the training pipeline, the overall input resolution was constrained by a global upper bound of 6,422,528 pixels (MAX_PIXELS), preventing overflow during multi-GPU parallel training and ensuring consistent input scaling across image and video modalities.
4.2 Main Results
Table 4: Comparison on the Chameleon [aide_yan2025a]. Accuracy (%) of different detectors (rows) in detecting real and fake images. For each training dataset, the first row is overall accuracy, the second row is “fake/real” class accuracy.
CNNSpot FreDect Fusing GramNet LNP UnivFD DIRE PatchCraft NPR AIDE Ivy-xDetector
60.89 57.22 57.09 60.95 58.52 60.42 59.71 56.32 58.13 65.77 73.17
9.86/99.25 0.89/99.55 0.02/99.98 4.76/99.66 7.72/96.70 85.52/41.56 11.86/95.67 3.07/96.35 2.43/100.00 26.80/95.06 67.78/77.22
@@ 4.2 Main Results
Table 5: Comparison on the Genimage [genimage_2023]. Accuracy (%) of different detectors (rows) in detecting real and fake images from different generators (columns). The best result and the second-best result are marked in bold and underline, respectively. Results of other methods are reported from [aide_yan2025a].
@@
CNNSpot [cnnspot_wang2020cnn] 52.80 96.30 95.90 50.10 39.80 78.60 53.40 46.80 64.21
F3Net [qian2020thinking] 50.10 99.90 99.90 49.90 50.00 99.90 49.90 49.90 68.69
DIRE [dire_wang2023dire] 60.20 99.90 99.80 50.90 55.00 99.20 50.10 50.20 70.66
GenDet [zhu2023gendet] 89.60 96.10 96.10 58.00 78.40 92.80 66.50 75.00 81.56
PatchCraft [aigcdetectionbenchmark_zhong2023patchcraft] 79.00 89.50 89.30 77.30 78.40 89.30 83.70 72.40 82.30
FRIDA [bonechi2025thisfakedetectionsource] 85.80 85.00 85.50 84.00 87.30 83.90 86.50 84.80 85.30
AIDE [aide_yan2025a] 79.38 99.74 99.76 78.54 91.82 98.65 80.26 66.89 86.88
DRCT [drct] 91.50 95.00 94.40 79.40 89.20 94.70 90.00 81.70 89.50
Effort [effort] 82.40 99.80 99.80 78.70 93.30 97.40 91.70 77.60 91.10
ThinkFake-7B [huang2025thinkfakereasoningmultimodallarge] 92.50 93.10 95.30 73.10 87.40 93.60 66.20 70.80 84.00
Ivy-xDetector 96.41 97.07 97.14 95.59 95.95 96.92 96.03 95.22 96.32
Table 6: Comparisons to the GenVideo [demamba]. F1 score (F1), recall score (R) and average precision (AP) on the many-to-many generalization task. “Demamba-XCLIP-FT” is abbreviated as “Demamba”. Results of other methods are reported from [demamba].
@@
Ivy-xDetector R 0.7857 0.9371 0.9507 0.9443 0.9550 0.9643 0.9968 0.9857 0.8943 0.9461 0.9528
F1 0.8800 0.9676 0.9747 0.9713 0.9770 0.9818 0.9984 0.9928 0.9442 0.9723 0.9526
We perform extensive experiments to assess both detection and explanation capabilities. In particular, the proposed method is evaluated on the classification (real/fake) tasks and reasoning tasks for both image and video content using the proposed unified framework. For the classification task, we test our model on both image and video content to detect the synthesic content.
From Table 3, our method demonstrates robust detection and explanation capabilities across both image and video tasks. Overall, closed-source models maintain a clear advantage in classification accuracy and subjective explanation quality. Among them, Gemini-2.5-flash achieves the best performance on both automatic metrics and GPT-assisted evaluation (e.g., reaching Acc/F1 of 0.812/0.812 on video detection and an average explanation score above 4.0), reflecting strong overall detection and reasoning ability. In contrast, GPT-4o maintains relatively high classification accuracy.
For open-source models, most models in the 3B–7B parameter range exhibit relatively low scores on explanation dimensions (<3.5<3.5), indicating limitations in generating reasoning chains and covering multi-dimensional details. Notably, Qwen2.5-VL-3B shows relatively stable performance in explanation relevance and detail (Det./Exp. >3.2>3.2), suggesting that lightweight models still hold potential under specific designs.
Notably, Gemini-2.5-Flash achieves the highest scores across all explanation-related dimensions, with an average of 4.12, surpassing all other models. This advantage can be attributed to the fact that part of our training data was distilled from Gemini-2.5-Pro, a powerful multimodal model capable of handling complex reasoning and explanation tasks. In contrast, our Ivy-xDetector ranks second with an average explanation score of 3.85, demonstrating interpretability comparable to Gemini-2.5-Flash while achieving substantially higher detection accuracy (Acc/F1 = 0.864/0.864) than any other open-source model. Compared with Gemini-2.5-Flash (Acc/F1 = 0.780/0.776), Ivy-xDetector improves accuracy by over 10 percentage points, confirming its robustness in both synthetic image detection and interpretive reasoning tasks.
4.3 Generalization Evaluation
Evaluation on Classification Benchmarks.
We conduct evaluations across multiple public leaderboards, including the image-based GenImage 5 and the video-based GenVideo benchmark 6. Performance results of other competing detectors are taken from AIDE [aide_yan2025a] and Demamba [demamba]. Our method achieves superior performance on all these leaderboards, demonstrating superior detection accuracy and robust cross-modal consistency.
Evaluation on Unseen Benchmarks.
Beyond the datasets aligned with our training sets, we further evaluate the generalization ability of Ivy-xDetector on the Chameleon benchmark[aide_yan2025a], which lies outside the training data distribution, as shown in Table 4. The results confirm that our approach preserves strong generalization capability across unseen generative models and diverse data domains.
Compared to the original leaderboard SOTA method AIDE, our model achieves an overall accuracy of 73.17%. Notably, while the accuracy for the real class decreases from 95.06% to 77.22%, the accuracy for the fake class substantially increases from 26.80% to 67.78%. This suggests that AIDE tends to overfit the real category, whereas our model achieves a more balanced and robust performance across both real and synthetic content.
Ablation Study.
Table 7: Ablation study on different SFT and RL training settings. Base Model denotes the base model Qwen2.5-VL-3B. Evaluation metric: ACC / F1.
Model Name Training Stage Chameleon (Image) GenVideo (Video) Ivy-Fake Test
SFT RL
Base Model 59.05 / 53.30 67.70 / 80.74 65.10 / 61.80
- 40.31 / 37.99 91.61 / 91.55 73.91 / 73.83
- 72.72 / 70.79 78.19 / 77.62 77.00 / 76.66
- 73.17 / 73.14 95.29 / 95.29 86.40 / 86.30
To validate the effectiveness of our training framework, we design four ablation variants, as shown in Table 7. As shown, applying SFT alone results in only a slight performance drop on the out-of-domain benchmark Chameleon, while achieving consistent improvements across all other datasets. This suggests that the SFT stage primarily enhances the model’s fine-grained sensitivity to synthetic artifacts. In contrast, applying RL alone yields modest but consistent gains in classification accuracy across benchmarks. The combination of SFT and RL achieves the best balance between accuracy and interpretability.
4.4 Case Study
As illustrated in Figure 5, Ivy-xDetector demonstrates a rich single frame / video understanding across multiple dimensions, effectively capturing subtle differences in generated content. To further highlight its advantages over other models, we provide additional comparison cases with GPT-4o, Gemini-2.5, InternVL3, and Qwen2.5-VL in Appendix 14.